# Modularity aided consistent attributed graph clustering via coarsening

### v1

@@ 5 Integrating with GNNs
Our optimization framework integrates seamlessly with Graph Neural Networks (GNNs) by incorporating the optimization objective equation 4 into the loss function. This integration can be minimized using gradient descent. We demonstrate the effectiveness of this approach on several popular GNN architectures, including Graph Convolutional Networks (GCNs) (Kipf & Welling, 2016b), Variational Graph Auto-Encoders (VGAEs) (Kipf & Welling, 2016a), and a variant known as Gaussian Mixture Model VGAE (GMM-VGAE) (Hui et al., 2020).
We iteratively learn the matrix $C$ using gradient descent. From equation 1, we update $X_{C}$ using the relation $X_{C}=C^{\dagger}X$. It is important to note that the loss term $CX_{C}-X$ will not necessarily be zero. This arises because we use a "soft" version of $C$ in the loss function to enable gradient flow, while a "hard" version of $C$ is used in the update step. This method ensures that $C$ naturally becomes harder and exhibits a higher prediction probability.
@@ 6.3 Attributed Graph Clustering
Cora CiteSeer PubMed Method ACC $\uparrow$ NMI $\uparrow$ ARI $\uparrow$ ACC $\uparrow$ NMI $\uparrow$ ARI $\uparrow$ ACC $\uparrow$ NMI $\uparrow$ ARI $\uparrow$ K-means 34.7 16.7 25.4 38.5 17.1 30.5 57.3 29.1 57.4 Spectral Clustering 34.2 19.5 30.2 25.9 11.8 29.5 39.7 3.5 52.0 DeepWalk (Perozzi et al., 2014a) 46.7 31.8 38.1 36.2 9.7 26.7 61.9 16.7 47.1 Louvain (Blondel et al., 2008b) 52.4 42.7 24.0 49.9 24.7 9.2 30.4 20.0 10.3 GAE [NeurIPS’16] (Kipf & Welling, 2016a) 61.3 44.4 38.1 48.2 22.7 19.2 63.2 24.9 24.6 DGI [ICLR’19] (Veličković et al., 2019) 71.3 56.4 51.1 68.8 44.4 45.0 53.3 18.1 16.6 GIC [PAKDD’21] (Mavromatis & Karypis, 2021) 72.5 53.7 50.8 69.6 45.3 46.5 67.3 31.9 29.1 DAEGC [IJCAI’19] (Wang et al., 2019) 70.4 52.8 49.6 67.2 39.7 41.0 67.1 26.6 27.8 GALA [ICCV’19] (Park et al., 2019) 74.5 57.6 53.1 69.3 44.1 44.6 69.3 32.1 32.1 AGE [KDD’20] (Cui et al., 2020) 73.5 57.5 50.0 69.7 44.9 34.1 71.1 31.6 33.4 DCRN [AAAI’22] (Liu et al., 2022) 61.9 45.1 33.1 70.8 45.8 47.6 69.8 32.2 31.4 FGC [JMLR’23] (Kumar et al., 2023) 53.8 23.2 20.5 54.2 31.1 28.2 67.1 26.6 27.8 Q-MAGC (Ours) 65.8 51.8 42.0 65.9 40.8 40.1 66.7 32.8 27.9 Q-GCN (Ours) 71.6 58.3 53.6 71.5 47.0 49.1 64.1 32.1 26.5 SCGC [IEEE TNNLS’23] (Liu et al., 2023a) 73.8 56.1 51.7 71.0 45.2 46.2 - - - MVGRL [ICML’20] (Hassani & Khasahmadi, 2020) 73.2 56.2 51.9 68.1 43.2 43.4 69.3 34.4 32.3 VGAE [NeurIPS’16] (Kipf & Welling, 2016a) 64.7 43.4 37.5 51.9 24.9 23.8 69.6 28.6 31.7 ARGA [IJCAI’18] (Pan et al., 2018) 64.0 35.2 61.9 57.3 34.1 54.6 59.1 23.2 29.1 ARVGA [IJCAI’18] (Pan et al., 2018) 63.8 37.4 62.7 54.4 24.5 52.9 58.2 20.6 22.5 R-VGAE [IEEE TKDE’22] (Mrabah et al., 2022) 71.3 49.8 48.0 44.9 19.9 12.5 69.2 30.3 30.9 Q-VGAE (Ours) 72.7 58.6 49.6 66.1 47.4 50.2 64.3 32.6 28.0 VGAECD-OPT [Entropy’20] (Choong et al., 2020) 27.2 37.3 22.0 51.8 25.1 15.5 32.2 25.0 26.1 Mod-Aware VGAE [NN’22] (Salha-Galvan et al., 2022) 67.1 52.4 44.8 51.8 25.1 15.5 - 30.0 29.1 GMM-VGAE [AAAI’20] (Hui et al., 2020) 71.9 53.3 48.2 67.5 40.7 42.4 71.1 29.9 33.0 R-GMM-VGAE [IEEE TKDE’22] (Mrabah et al., 2022) 76.7 57.3 57.9 68.9 42.0 43.9 74.0 33.4 37.9 Q-GMM-VGAE (Ours) 76.2 58.7 56.3 72.7 47.4 48.8 69.0 34.8 34.0
@@ 6.5 Ablation Studies
(a) Comparison of all methods on non-attributed datasets using degree.
(a) Comparison of running times of methods.
Note that the scale for PubMed is in minutes (right axis), whereas for Cora and CiteSeer is in seconds.

### v2

@@ 5 Integrating with GNNs
Our optimization framework integrates seamlessly with Graph Neural Networks (GNNs) by incorporating the objective equation 4 into the loss function. This integration can be minimized using gradient descent. We demonstrate the effectiveness of this approach on several popular GNN architectures, including Graph Convolutional Networks (GCNs) (Kipf & Welling, 2016b), Variational Graph Auto-Encoders (VGAEs) (Kipf & Welling, 2016a), and a variant known as Gaussian Mixture Model VGAE (GMM-VGAE) (Hui et al., 2020).
We iteratively learn the matrix $C$ using gradient descent. From equation 1, we update $X_{C}$ using the relation $X_{C}=C^{\dagger}X$. It is important to note that the loss term $CX_{C}-X$ will not necessarily be zero. This arises because we use a "soft" version of $C$ ($C_{i,j}$ is the probability $\in[0,1]$) in the loss function to enable gradient flow, while a "hard" version of $C$ ($C_{i,j}$ is the binary assignment $\in\{0,1\}$) is used in the update step. This method ensures that $C$ naturally becomes harder and exhibits a higher prediction probability.
@@ 6.3 Attributed Graph Clustering
Cora CiteSeer PubMed Method ACC $\uparrow$ NMI $\uparrow$ ARI $\uparrow$ ACC $\uparrow$ NMI $\uparrow$ ARI $\uparrow$ ACC $\uparrow$ NMI $\uparrow$ ARI $\uparrow$ K-means 34.7 16.7 25.4 38.5 17.1 30.5 57.3 29.1 57.4 Spectral Clustering 34.2 19.5 30.2 25.9 11.8 29.5 39.7 3.5 52.0 DeepWalk (Perozzi et al., 2014a) 46.7 31.8 38.1 36.2 9.7 26.7 61.9 16.7 47.1 Louvain (Louvain_Blondel_2008) 52.4 42.7 24.0 49.9 24.7 9.2 30.4 20.0 10.3 GAE [NeurIPS’16] (Kipf & Welling, 2016a) 61.3 44.4 38.1 48.2 22.7 19.2 63.2 24.9 24.6 DGI [ICLR’19] (Veličković et al., 2019) 71.3 56.4 51.1 68.8 44.4 45.0 53.3 18.1 16.6 GIC [PAKDD’21] (Mavromatis & Karypis, 2021) 72.5 53.7 50.8 69.6 45.3 46.5 67.3 31.9 29.1 DAEGC [IJCAI’19] (Wang et al., 2019) 70.4 52.8 49.6 67.2 39.7 41.0 67.1 26.6 27.8 GALA [ICCV’19] (Park et al., 2019) 74.5 57.6 53.1 69.3 44.1 44.6 69.3 32.1 32.1 AGE [KDD’20] (Cui et al., 2020) 73.5 57.5 50.0 69.7 44.9 34.1 71.1 31.6 33.4 DCRN [AAAI’22] (Liu et al., 2022) 61.9 45.1 33.1 70.8 45.8 47.6 69.8 32.2 31.4 FGC [JMLR’23] (Kumar et al., 2023) 53.8 23.2 20.5 54.2 31.1 28.2 67.1 26.6 27.8 Q-MAGC (Ours) 65.8 51.8 42.0 65.9 40.8 40.1 66.7 32.8 27.9 Q-GCN (Ours) 71.6 58.3 53.6 71.5 47.0 49.1 64.1 32.1 26.5 SCGC [IEEE TNNLS’23] (Liu et al., 2023a) 73.8 56.1 51.7 71.0 45.2 46.2 - - - MVGRL [ICML’20] (Hassani & Khasahmadi, 2020) 73.2 56.2 51.9 68.1 43.2 43.4 69.3 34.4 32.3 VGAE [NeurIPS’16] (Kipf & Welling, 2016a) 64.7 43.4 37.5 51.9 24.9 23.8 69.6 28.6 31.7 ARGA [IJCAI’18] (Pan et al., 2018) 64.0 35.2 61.9 57.3 34.1 54.6 59.1 23.2 29.1 ARVGA [IJCAI’18] (Pan et al., 2018) 63.8 37.4 62.7 54.4 24.5 52.9 58.2 20.6 22.5 R-VGAE [IEEE TKDE’22] (Mrabah et al., 2022) 71.3 49.8 48.0 44.9 19.9 12.5 69.2 30.3 30.9 Q-VGAE (Ours) 72.7 58.6 49.6 66.1 47.4 50.2 64.3 32.6 28.0 VGAECD-OPT [Entropy’20] (Choong et al., 2020) 27.2 37.3 22.0 51.8 25.1 15.5 32.2 25.0 26.1 Mod-Aware VGAE [NN’22] (Salha-Galvan et al., 2022) 67.1 52.4 44.8 51.8 25.1 15.5 - 30.0 29.1 GMM-VGAE [AAAI’20] (Hui et al., 2020) 71.9 53.3 48.2 67.5 40.7 42.4 71.1 29.9 33.0 R-GMM-VGAE [IEEE TKDE’22] (Mrabah et al., 2022) 76.7 57.3 57.9 68.9 42.0 43.9 74.0 33.4 37.9 Q-GMM-VGAE (Ours) 76.2 58.7 56.3 72.7 47.4 48.8 69.0 34.8 34.0
@@ 6.5 Ablation Studies
(a) Comparison of all methods on non-attributed datasets.
(a) Comparison of running times of methods. scale for PubMed (right axis): mins, scale for others (left axis) : secs.
@@ Appendix I Attributed SBM theory and results
We selected this widely used and studied model for our analysis primarily because it considers a degree parameter for all nodes, resembling a key characteristic of real-world graphs. Therefore, the DC-SBM is closer to real-world graphs than the simpler SBM while still being analyzable, making it relevant here.